# Learning to Generate Object Interactions with Physics-Guided Video Diffusion

### v1

@@ Abstract
Recent models for video generation have achieved remarkable progress and are now deployed in film, social media production, and advertising. Beyond their creative potential, such models also hold promise as world simulators for robotics and embodied decision making. Despite strong advances, however, current approaches still struggle to generate physically plausible object interactions and lack physics-grounded control mechanisms. To address this limitation, we introduce KineMask, an approach for physics-guided video generation that enables realistic rigid body control, interactions, and effects. Given a single image and a specified object velocity, our method generates videos with inferred motions and future object interactions. We propose a two-stage training strategy that gradually removes future motion supervision via object masks. Using this strategy we train video diffusion models (VDMs) on synthetic scenes of simple interactions and demonstrate significant improvements of object interactions in real scenes. Furthermore, KineMask integrates low-level motion control with high-level textual conditioning via predictive scene descriptions, leading to effective support for synthesis of complex dynamical phenomena. Extensive experiments show that KineMask achieves strong improvements over recent models of comparable size. Ablation studies further highlight the complementary roles of low- and high-level conditioning in VDMs. Our code, model, and data will be made publicly available.
Project Page: https://daromog.github.io/KineMask/
@@ 1 Introduction
Recent years have seen substantial advances in video generation, with Video Diffusion Models (VDMs) emerging as a leading paradigm for high-resolution, temporally consistent synthesis (Ho et al., 2020; Blattmann et al., 2023b; Yang et al., 2025; Kong et al., 2024). This progress has elevated visual quality and enabled early commercial use in creative performances (Miller, 2023), experimental filmmaking (Chayka, 2023), and advertising (Roth, 2025). Beyond content creation, VDMs have also been explored as world models (Ha & Schmidhuber, 2018), capable of anticipating real-world interactions and supporting robotics and embodied decision making (Agarwal et al., 2025; Alonso et al., 2024; Ding et al., 2024). Realizing this vision, however, requires strict physical plausibility and control, since small deviations from realistic physics can accumulate into large errors in predicted dynamics. Yet current VDMs struggle to capture fundamental traits (Motamed et al., 2025; Kang et al., 2025) such as object permanence and causal interactions. Hence, methods for physically accurate video generation are a key step toward establishing VDMs as reliable world models.
Figure 1: KineMask results. We enable object-based control with a novel training strategy. Paired with synthetic data constructed for the task, KineMask enables pretrained diffusion models to synthesize realistic object interactions in real-world input scenes.
@@
Recent frameworks such as PhysGen (Liu et al., 2024b) and WonderPlay (Li et al., 2025c) integrate physics-based simulators into data-driven video generation, but rely on explicit scene reconstruction, a challenging task on its own. Around the time of this work, Force Prompting (Gillman et al., 2025) introduced physics-guided controls into VDMs using simulated training data. While promising, it remains limited in handling rigid-body interactions, often producing unrealistic dynamics, and distorted shapes. Motivated by these limitations, we propose an approach that enables video diffusion models to generate physically realistic object interactions, as a core requirement for robotics and advanced video generation. We focus on two central questions: (1) Can a video diffusion model generate realistic interactions between objects given initial dynamic conditions, and (2) how do data and textual conditioning influence the emergence of causal physical effects in generated videos?
@@
To address these questions, we introduce KineMask, a physics-guided framework for generating object interactions and effects in complex scenes (Figure 1). KineMask provides low-level kinematic control over parameters such as object direction and speed, while inferring object interactions directly within the VDM. KineMask is trained on simulator-rendered videos that capture not only physically valid dynamics but also explicit object interactions, paired with textual descriptions of the underlying events. We further leverage object masks as guidance, which enables refined control over motion trajectories and improves the model’s understanding of object shapes. Beyond low-level control, KineMask integrates high-level prompt conditioning through textual descriptions of future scene dynamics. At inference, it predicts dynamics from an input image and enables the generation of complex effects such as glass shattering or liquid spilling. Extensive experiments demonstrate that KineMask not only introduces new control capabilities but also outperforms state-of-the-art models of comparable size, while ablation studies highlight the importance of both the proposed training strategy and the integration of low- and high-level controls.
@@
  • We train KineMask on a synthetic dataset of dynamic scenes with simple object interactions and demonstrate generalization of our model to complex interactions in real scenes.
  • We combine low-level motion control with high-level text conditioning, yielding significant gains over comparable models in realistic video synthesis.
@@ 2 Related works
Early video diffusion directly extended image generators by inserting temporal layers into denoising U-Nets (Agarwal et al., 2022; Blattmann et al., 2023a; Guo et al., 2024b; Wang et al., 2023; 2025b; Chen et al., 2023; Bar-Tal et al., 2024). Later, video synthesis improved through the usage of Diffusion Transformers (DiTs) (Peebles & Xie, 2023), inheriting scaling properties from native transformer-based architectures. The use of DiTs allowed for higher-resolution videos and stronger visual quality (Yang et al., 2025; Kong et al., 2024; Wan et al., 2025; HaCohen et al., 2024). We build KineMask on CogVideoX (Yang et al., 2025), benefiting from the advantages of DiTs. Beyond creative purposes, diffusion-based video generation is now used for world modeling, by synthesizing the possible outcomes of actions in an environment. (Alonso et al., 2024; Valevski et al., 2025) first proposed diffusion-based world models on restricted scenarios. Some like GAIA (Hu et al., 2023; Russell et al., 2025) train at scale on specific domains such as autonomous driving, while Cosmos (Agarwal et al., 2025) uses large-scale data to handle heterogeneous domains. These models still suffer from limited realism in generating object interactions, motivating our study.
@@
Significant efforts have been made to extend control on generated videos beyond textual conditioning. On images, ControlNet (Zhang et al., 2023) and similar approaches (Mou et al., 2024; Zhao et al., 2023) proposed plug-and-play control trainable components for dense conditioning such as edges, depth, or human poses. In videos, VideoControlNet (Hu & Xu, 2024) propagated conditions frame by frame using optical flow, while others focused on pose-driven human video generation (Zhang et al., 2024b) or keyframe-to-video propagation (Li et al., 2024a). More recent works such as SparseCtrl (Guo et al., 2024a) uses just a few keyframes for sketch, depth, or image conditioning, ignoring motion conditions. On motion, DragDiffusion (Shi et al., 2024) and MotionCtrl (Li et al., 2024b) allow interactive editing and trajectory-aware conditioningm. Motion Prompting (Geng et al., 2025) introduces point-track motion prompts, Tora (Zhang et al., 2025) designs trajectory-oriented diffusion transformers. All these require pre-defined trajectories. Some drive motion by directly guiding attentions (Pondaven et al., 2025) or noise (Burgert et al., 2025) using reference videos. Beyond single modalities, Cosmos-Transfer (Abu Alhaija et al., 2025) demonstrates adaptive multimodal conditioning. To our knowledge, KineMask is the first approach designed to generate object interaction by controlling initial object velocity.
@@
The interactions between physical understanding and video generation is a growing research field. A first line of work integrates physical simulations with learning-based techniques. In particular, PhysDreamer (Zhang et al., 2024a) generates oscillatory motions on 3D Gaussians, while DreamPhysics learns physical properties of dynamic 3D Gaussians with video diffusion priors (Huang et al., 2025). On video generation, WonderPlay (Li et al., 2025c) bridges physics solvers and generative video models to synthesize dynamic 3D scenes across diverse physical phenomena. PhysGen (Liu et al., 2024b) applies a similar approach using only 2D information for rigid body interactions. However, the use of a simulator requires significant engineering efforts and limits the flexibility of models. Differently, C-Drag (Li et al., 2025b) uses a LLM for inferring causal motion on output videos, also requiring tracking-based control. Using purely diffusion models, InterDyn (Akkerman et al., 2025) explores the capability of video models to render realistic object dynamics. However, it exploits frame-wise masks of controlling elements, which are typically unavailable at test time. Similarly to us, Li et al. (2025a) explores physics post-training but focuses on gravity effects. The concurrent work Force Prompting (Gillman et al., 2025) explores similar ideas to ours but does not consider object interactions and deploys simpler motion control. We experimentally compare to Force Prompting and demonstrate improved performance.
@@ 3 KineMask
KineMask is designed to synthesize realistic interactions among objects given the image of an initial scene and initial object velocity encoded by the object mask, see Figure 2. Below we introduce preliminaries in Section 3.1, we then describe our conditioning mechanism in Section 3.2 and outline the training and inference procedure in Section 3.3.
@@ 3.1 Preliminaries
$\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c,\mathbf{y}),$ (1)
@@
$\mathcal{L}_{\text{diff}}(\theta;\mathbf{x}_{0},\mathbf{y},c,t)=D_{\mathrm{KL}}\!\Big(p(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})\;\big\|\;p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c,\mathbf{y})\Big).$ (2)
@@
To allow additional guidance, a ControlNet (Zhang et al., 2023) branch $\psi_{\phi}$, parameterized by $\phi$, can encode an arbitrary dense control signal $\mathbf{u}\in\mathbb{R}^{F\times H\times W\times D}$ driving the output generation. For more details, please refer to Zhang et al. (2023). The denoising step then becomes:
@@
When training the ControlNet, the parameters of the backbone model $\theta$ are kept frozen, while only the control branch $\phi$ is optimized. The corresponding loss is
$\mathcal{L}_{\text{ctrl}}(\phi;\mathbf{x}_{0},\mathbf{y},\mathbf{u},c,t)=D_{\mathrm{KL}}\!\Big(q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})\;\big\|\;p_{\theta}\!\bigl(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c,\mathbf{y},\psi_{\phi}(\mathbf{u})\bigr)\Big).$ (4)
@@ 3.2 Enabling motion control
We now want to enable object-wise motion control for KineMask. Specifically, our goal is to move an object in an input scene $\mathbf{y}$ with controlled direction and velocity, allowing us to study the effects of object interactions in the videos generated with diffusion models. To do so, we assume access to a dataset $\mathcal{D}$ of captioned videos depicting objects in motion. Let $f\in\{1,\dots,F\}$ denote the frame index. For each frame $f$ and object in the scene, we are given a mask $\mathbf{m}_{f}\in\mathbb{R}^{H\times W\times 3}$ aligned with the image resolution. The three channels encode the instantaneous velocity vector, with the red, green, and blue channels corresponding to motion along the $x$-, $y$-, and $z$-axes, respectively, in the pixels defined by a segmentation mask of the object. In this way, $\mathcal{D}$ provides not only spatial information about object locations, but also explicit ground-truth dynamics in three dimensions. The velocity masks are then aggregated into a tensor $\mathbf{m}\in\mathbb{R}^{F\times H\times W\times 3}$ and used to condition $\psi_{\phi}$. Similarily to Akkerman et al. (2025), we annotate in this way only the velocity of the objects moving in the first frame of the rendered video, leaving blank the mask for objects potentially moved by interactions. We visualize our strategy in Figure 2 (top). This enforces the model to synthesize interactions without explicitly relying on pixel control information. We can then train a first-stage KineMask ControlNet $\phi^{\prime}$ by solving
@@
Figure 2: KineMask pipeline. We encode our low-level control signal as a mask encoding the instantaneous velocity of the moving objects for each frame, to train a ControlNet (left) in two stages using Blender-generated videos of objects in motion. In the first one, we train with all frames, whereas in the second one, we randomly drop part of the final frames. We also provide a high-level textual control extracted by a VLM. At inference (right), we construct the low-level conditioning with SAM and use GPT to infer high-level outcomes of object motion from a single frame.
@@
KineMask $\phi^{\prime}$ enables motion control given motion masks $\mathbf{m}$ provided for all the frames of a video. While such a setup simplifies training, it does not correspond to our desired scenario of video generation conditioned only by the object motion at the first video frame. Towards this goal, we propose a mask dropout strategy, erasing the last part of the velocity masks in $\mathbf{m}$ at training time, as shown in Figure 2 (bottom). Formally, we define a truncated mask tensor
$\mathbf{m}_{\odot}=\{\mathbf{m}_{\odot,f}=\mathbf{m}_{f}\;\;\text{if }f\leq f^{*},\;\;\mathbf{0}\;\;\text{otherwise}\}_{f=0}^{F}.$ (6)
where $f^{*}$ denotes the cutoff frame index corresponding to the dropout ratio. Thus, only the first frames contain velocity supervision, while the remainder of the sequence is set to zero. We then train a second-stage KineMask $\phi^{\prime\prime}$ by finetuning $\phi^{\prime}$ with this strategy, solving
@@ 3.3 Data
At inference time, we assume as input an unseen image $\mathbf{y}$. An object mask can be easily obtained for the target object e.g., using SAM2 (Ravi et al., 2025), while the desired object velocity at the first frame is assumed to be provided by the user. We use this information to construct $\mathbf{m}_{\odot}$. We also prompt GPT-5 (OpenAI, 2025) for a description $c_{\text{infer}}$ of the effects on the scene if the object starts moving in the direction indicated by the user. The full prompt is in Appendix A.2. Combining those with random noise $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$, we construct the input tuple $\{\mathbf{x}_{T},\mathbf{y}_{\text{input}},\mathbf{m}_{\odot},c_{\text{infer}}\}$ compatible with our VDM equipped with $\phi^{\prime\prime}$. Our inference pipeline is illustrated in Figure 2 (right).

### v2

# Learning to Generate Rigid Body Interactions with Video Diffusion Models

@@ Abstract
Recent video generation models have achieved remarkable progress and are now deployed in film, social media production, and advertising. Beyond their creative potential, such models also hold promise as world simulators for robotics and embodied decision making. Despite strong advances, however, current approaches still struggle to generate physically plausible object interactions and lack object-level control mechanisms. To address these limitations, we introduce KineMask, an approach for video generation that enables realistic rigid body control, interactions, and effects. Given a single image and a specified object velocity, our method generates videos with inferred motions and future object interactions. We propose a two-stage training strategy that gradually removes future motion supervision via object masks. Using this strategy we train video diffusion models (VDMs) on synthetic scenes of simple interactions and demonstrate significant improvements of rigid body interactions in real scenes. Furthermore, KineMask integrates low-level motion control with high-level textual conditioning via predicted scene descriptions, leading to support for synthesis of complex dynamical phenomena. Our experiments show that KineMask achieves strong improvements over recent models of comparable size. Ablation studies further highlight the complementary roles of low- and high-level conditioning in VDMs. Project Page: https://daromog.github.io/KineMask/
Figure 1: KineMask results. We enable object-based control with a novel training strategy. Paired with synthetic data constructed for the task, KineMask enables pretrained diffusion models to synthesize realistic rigid body interactions in real-world input scenes.
@@ 1 Introduction
Recent years have seen substantial advances in video generation, with Video Diffusion Models (VDMs) emerging as a leading paradigm for high-resolution, temporally consistent synthesis [25, 10, 61, 31]. This progress has elevated visual quality and enabled early commercial use in creative performances [40], experimental filmmaking [12], and advertising [47]. Beyond content creation, VDMs have also been explored as world models [23], capable of anticipating real-world interactions and supporting robotics and embodied decision making [3, 5, 18]. Realizing this vision, however, requires strict physical plausibility and control, since small deviations from realistic physics can accumulate into large errors in predicted dynamics. Yet, current VDMs struggle to capture fundamental traits [41, 30] such as object permanence, collisions and causal interactions, even for large-scale models like Veo-3 [60]. Hence, methods for physically accurate video generation are a key step toward establishing VDMs as reliable world models.
@@
Recent frameworks such as WonderPlay [37], PhysGen  [39] and PhysGen3D [13] integrate physics-based simulators into data-driven video generation, but rely on explicit scene reconstruction, a challenging task on its own. Differently, drag-based approaches [50, 35, 66, 19, 68, 32, 15] can move objects, but often requiring pre-defined target points, thus preventing the inference of causal effects of motion from starting conditions only. Around the time of this work, Force Prompting [20] introduced physics-guided controls into VDMs using simulated training data. While promising, it offers limited fine-grained control and object interactions realism, often producing unrealistic dynamics, and distorted shapes, showing a lack of understanding of causality. Motivated by these limitations, we propose an approach enabling video diffusion models to generate physically realistic object interactions, as a core requirement for robotics and video generation. Our approach aims to answer two central questions for world models: (1) Can a video diffusion model generate realistic interactions between rigid bodies given initial dynamic conditions, and (2) how do data and textual conditioning influence the emergence of causal physical effects in generated videos?
@@
To address these questions, we introduce KineMask, a framework for generating accurate object interactions and effects in complex scenes (Figure 1). KineMask provides low-level kinematic control over parameters such as object direction and speed, while inferring object interactions directly within the VDM. KineMask is trained on simulator-rendered videos that capture not only physically valid dynamics but also explicit object interactions, paired with textual descriptions of the underlying events. We leverage object masks as guidance, enabling single-frame motion control and improving the model’s understanding of shapes. Beyond low-level control, KineMask integrates high-level prompt conditioning through textual descriptions of future scene dynamics. At inference, it predicts dynamics from an input image and enables the generation of complex effects, such as liquid spilling (Figure 1). Extensive experiments show that KineMask not only introduces new control capabilities but also outperforms state-of-the-art models of comparable size, while ablation studies highlight the importance of both the proposed training strategy and the integration of low- and high-level controls.
@@
  • We train KineMask on a synthetic dataset of dynamic scenes with simple object interactions and demonstrate generalization to complex interactions in real scenes.
  • We combine low-level motion control with high-level text conditioning, yielding significant gains over comparable models in video synthesis of rigid body interactions.
@@ 2 Related works
Early video diffusion directly extended image generators by inserting temporal layers into denoising U-Nets [2, 9, 22, 57, 59, 14, 7]. Later, video synthesis improved through the usage of Diffusion Transformers (DiTs) [44], inheriting scaling properties from native transformer-based architectures. The use of DiTs allowed for higher-resolution videos and stronger visual quality [61, 31, 55, 24]. We build KineMask on CogVideoX [61], benefiting from the advantages of DiTs. Beyond creative purposes, diffusion-based video generation is now used for world modeling, by synthesizing the possible outcomes of actions in an environment. [5, 54] first proposed diffusion-based world models on restricted scenarios. Some like GAIA [26, 48] train at scale on specific domains such as autonomous driving, while Cosmos [3] uses large-scale data to handle heterogeneous domains. These models still suffer from limited realism in generating object interactions, motivating our study.
@@
Significant efforts have been made to extend control on generated videos beyond textual conditioning. On images, ControlNet [63] and similar approaches [42, 67] proposed plug-and-play control trainable components for dense conditioning such as edges, depth, or human poses. In videos, VideoControlNet [27] propagated conditions frame by frame using optical flow, while others focused on pose-driven human video generation [65] or keyframe-to-video propagation [34]. More recent works such as SparseCtrl [21] uses just a few keyframes for sketch, depth, or image conditioning, ignoring motion conditions. MotionI2V [49] first infers object motion with flow generation, then it renders a video using flow as control. A significant line of works allows motion control with dragging trajectories [50, 35, 66, 19, 68, 32, 17, 56, 62]. While effective, these require pre-defined trajectories, and are unable to infer motion causal effects from initial conditions. Some drive motion by directly guiding attentions [45] or noise [11] using reference videos. Beyond single modalities, Cosmos-Transfer [1] demonstrates adaptive multimodal conditioning. To our knowledge, KineMask is the first approach training VDMs to generate rigid body interactions by controlling initial object velocity only.
Figure 2: KineMask pipeline. We encode our low-level control signal as a mask encoding the velocity of the moving objects, to train a ControlNet (left) in two stages using Blender-generated videos of objects in motion. In the first one, we train with all frames, whereas in the second one, we randomly drop part of the final frames. We also provide a high-level textual control extracted by a VLM. At inference (right), we construct the low-level conditioning with SAM and use GPT to infer high-level outcomes of object motion from a single frame.
@@
The interactions between physical understanding and video generation is a growing research field. A first line of work integrates physical simulations with learning-based techniques. In particular, PhysDreamer [64] generates oscillatory motions on 3D Gaussians, while DreamPhysics learns physical properties of dynamic 3D Gaussians with video diffusion priors [28]. On video generation, WonderPlay [37] bridges physics solvers and generative video models to synthesize dynamic 3D scenes across diverse physical phenomena. PhysGen3D [13] and PhysGen [39] apply a similar approach using 3D and 2D information respectively for rigid body interactions. However, the use of a simulator requires significant engineering efforts and limits the flexibility of models. Differently, others [8, 16] map input kinematics to future frames, but requiring dataset-specific training. Instead, C-Drag [36] infers causal motion on output videos with LLMs, and combines it with tracking control to allow video generation. Using purely diffusion models, InterDyn [4] explores the capability of video models to render realistic object dynamics. However, it exploits frame-wise masks of controlling elements, which are typically unavailable at test time. Similarly to us, [33] explores physics post-training but focuses on gravity effects. Force Prompting [20] explores similar ideas to ours but does not consider object interactions and deploys simpler motion control. We compare to Force Prompting and demonstrate improved performance synthesizing object interactions.
@@ 3 KineMask
KineMask is designed to synthesize realistic interactions among rigid bodies given an input image and the object velocity encoded by an object mask, as in Figure 2. Below, we introduce preliminaries in Section 3.1. We then describe our conditioning mechanism in Section 3.2 and outline the training and inference procedure in Section 3.3.
@@ 3.1 Preliminaries
$\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c,\mathbf{y}),\text{\vskip-2.00749pt}$ (1)
@@
$\begin{split}&\mathcal{L}_{\text{diff}}(\theta;\mathbf{x}_{0},\mathbf{y},c,t)=D_{\mathrm{KL}}\!\big(\\ &~~~~~~~~~~~~~p(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})\;\big\|\;p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c,\mathbf{y})\big).\text{\vskip-2.00749pt}\end{split}$ (2)
@@
To allow additional guidance, a ControlNet [63] branch $\psi_{\phi}$, parameterized by $\phi$, can encode an arbitrary dense control signal $\mathbf{u}\in\mathbb{R}^{F\times H\times W\times D}$ driving the output generation. For more details, please refer to [63]. The denoising step then becomes:
@@
When training the ControlNet, the parameters of the backbone model $\theta$ are kept frozen, while only the control branch $\phi$ is optimized. The corresponding loss is:
$\begin{split}&\mathcal{L}_{\text{ctrl}}(\phi;\mathbf{x}_{0},\mathbf{y},\mathbf{u},c,t)=D_{\mathrm{KL}}\!\big(\\ &~~~~~~~~~~~~~q(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t},\mathbf{x}_{0})\,\big\|\,p_{\theta}(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t},c,\mathbf{y},\psi_{\phi}(\mathbf{u}))\big).\end{split}$ (4)
@@ 3.2 Enabling motion control
We now want to enable object-wise motion control for KineMask. Specifically, our goal is to move an object in an input scene $\mathbf{y}$ with controlled direction and velocity, allowing us to study the effects of object interactions in the videos generated with diffusion models. To do so, we assume access to a dataset $\mathcal{D}$ of captioned videos depicting objects in motion. Let $f\in\{1,\dots,F\}$ denote the frame index. For each frame $f$ and object in the scene, we are given a mask $\mathbf{m}_{f}\in\mathbb{R}^{H\times W\times 3}$ aligned with the image resolution. The three channels encode the instantaneous velocity vector, with the red, green, and blue channels corresponding to motion along the $x$-, $y$-, and $z$-axes, respectively, in the pixels defined by a segmentation mask of the object. In this way, $\mathcal{D}$ provides not only spatial information about object locations, but also explicit ground-truth dynamics. The velocity masks are then aggregated into a tensor $\mathbf{m}\in\mathbb{R}^{F\times H\times W\times 3}$ and used to condition $\psi_{\phi}$. Similarily to [4], we annotate in this way only the velocity of the objects moving in the first frame of the rendered video, leaving blank the mask for objects potentially moved by interactions. We visualize our strategy in Figure 2 (top). This enforces the model to synthesize interactions without explicitly relying on pixel control. We can then train a first-stage KineMask ControlNet $\phi^{\prime}$ by solving
@@
KineMask $\phi^{\prime}$ enables motion control given motion masks $\mathbf{m}$ provided for all the frames of a video. While such a setup simplifies training, it does not correspond to our desired scenario of video generation conditioned only by the object motion at the first video frame. Towards this goal, we use a mask dropout strategy, randomly erasing part of the velocity masks in $\mathbf{m}$ at training time, as shown in Figure 2 (bottom). Formally, we define a truncated mask tensor
$\mathbf{m}_{\odot}=\{\mathbf{m}_{\odot,f}=\mathbf{m}_{f}\;\;\text{if }f\in f^{*},\;\;\mathbf{0}\;\;\text{otherwise}\}_{f=0}^{F}.$ (6)
where $f^{*}$ denotes the masked frame indices, and depends on the dropout ratio. Thus, only some frames contain velocity supervision, while the remainder of the sequence is set to zero. We then train a second-stage KineMask $\phi^{\prime\prime}$ by finetuning $\phi^{\prime}$ with this strategy, solving
@@ 3.3 Data
We assume as input an unseen image $\mathbf{y}$. An object mask is obtained for the target object e.g., using SAM2 [46], while the desired object velocity at the first frame is assumed to be provided by the user. We use this information to construct $\mathbf{m}_{\odot}$. We also prompt GPT-5 [43] for a description $c_{\text{infer}}$ of the effects on the scene if the object starts moving in the direction indicated by the user. The full prompt is in Appendix A.2. Combining those with random noise $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$, we construct the input tuple $\{\mathbf{x}_{T},\mathbf{y}_{\text{input}},\mathbf{m}_{\odot},c_{\text{infer}}\}$ compatible with our VDM equipped with $\phi^{\prime\prime}$. Our inference pipeline is illustrated in Figure 2 (right). Please note that at inference the user simply draws an arrow, mapped to RGB masks automatically.
Figure 4: User study. We widely outperform baselines on motion fidelity, interaction quality, and overall physical consistency.